OpenStudy (anonymous):

3sqrtx-4=9 what is x? please help

OpenStudy (jdoe0001):

what do you think you'd do next? bearing in mind you'd want to isolate "x"

OpenStudy (anonymous):

uhm..you subtract 4 from 9? o.0 i really dont know

OpenStudy (jdoe0001):

yes :)

OpenStudy (jdoe0001):

so you'd be left with \(3\sqrt{x}=5\)

OpenStudy (jdoe0001):

then divide both by 3, then to get rid of the \(\sqrt{}\), raise both to 2 power

OpenStudy (anonymous):

yay (: thankyou soo much!

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

actually you ADD 4 to 9

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

so it would be 13 than? right?

OpenStudy (jdoe0001):

\(3\sqrt{x}-4(+4)=9(+4)\) and then the rest, divide by 3 and raise by 2power

OpenStudy (anonymous):

okay thankyou (:

OpenStudy (jdoe0001):

yw